Tutor,
I am not sure how to solve this problem. Could I get some help with this?
Find the altitude of the triangle.
13, 13, 10 are the sides.
Thank you for your help.

Answer by edjones(8007)

make the base 10
Then the altitude bisects the base and is perpendicular to it forming a right triangle with a base of 5 and hypotenuse of 13.
a^2+b^2=c^2
a^2+5^2=13^2
a^2+25=169
a^2=144
a=12
Ed